Shift report 20000715 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 15 2000 - 15:57:18 EDT

  • Next message: J. H. Lee: "Power Failure, and No beam is expected until Monday"

    ------------------------------------------------------------
    Date: Sat Jul 15 15:57:18 2000, Shift: 08:00-16:00
    Supervisor: BS, with: FF, TV, PC
    Log book#3 pages 17 - 21
    
    Shift Summary:
    ------------------------------------------------------------
    RHIC has had all sorts of problems, so no good beam 
    all shift. They managed some injection-testing until 11,
    but then they got a power dip and have been down since.
    
    No runs, no data... TPM2 and T2B had tripped when we came -
    this probably happened between 7.00 and 8.00.
    
    BrNtuple has been updated with new time-values for the BB.
    See logbook 3, page 19 for a before/after-plot.
    
    We can confirm MMs report of problems with TOFW - many
    slats are dead. We have not gotten hold of DB to discuss
    this yet.
    
    A question about TPC-voltages: TPM1 was moved from channel
    1.0 to 1.2, but we have a demand-voltage set on cannels 1.0,
    1.1 and 1.2. Are all three needed, or should one be 0?
    
    There has also been a lovely, wet storm. There is a large
    leak in hall just outside the bathroom and a few small ones
    elsewhere in 1002A and B - these should probably be
    fixed... We see no leaks in the counting-house, or in the IR
    (with the camera). (There is plastic beneath the roof in the
    FEH, so the water porbably can't get in there)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 15 2000 - 16:00:50 EDT